OpenStudy (anonymous):

Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. Express answers with positive exponents. (-27x^0y^-2/54x^-5y^-4)^-2

OpenStudy (anonymous):

(-27x^0y^-2/54x^-5y^-4)^-2 Exponent is negative, so flip the fraction! 54x^-5y^-4/-27x^0y^-2 54/-27 = -2 x^-5 = 1/x^5 y^-4/y^-2 = 1/y^2 -2 / x^5y^2 Now, square that. (-2/x^5y^2)^2 = 4/x^10y^4 4/x^10y^4